OpenStudy (anonymous):

the room is 2 meters longer than it is wide. its perimeter is 16 meters what is the width helppppppppppp!!!!!!!!!!!!!!

OpenStudy (anonymous):

l=length w=wide 2+w=l 16=2(2+w)+2w 16=4+2w+2w 16=4+4w -4 -4 12=4w divide both sides by 4 3=w now plug it in to see if it works 16=2(2+w)+2w 16=2(2+3)+2X3 16=2(5)+6 16=10+6 16=16

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np